# A pedagogical approach to the black hole information issue

### v3

@@ IV Time evolution of semiclassical black holes
In this section we show how the semiclassical properties of black holes discussed so far leads to the conclusion of information loss. We will show why this conclusion is in agreement with semiclassical arguments, and does not earn the adjective “paradoxical”. Additionally, we will present misconceptions surrounding some proposals that seek to change the conclusion of information loss.
@@ IV.1 Information loss
Although calculating general backreaction effects is not a trivial task for four-dimensional spacetimes, one expects that the main consequence of the Hawking radiation is to reduce the parameters of the black hole at a given time, i.e., its mass, angular momentum, and electric charge. Consequently, one would have an effective evaporation process. This qualitative description of a possible backreaction effect is corroborated by physical arguments. More precisely, the energy reduction is merely a consequence of the fact that the positive energy flux to $\mathscr{I}^{+}$ implies that there must exist a negative energy flux going into the black hole, while the reduction of the other parameters is due to the fact that the spectrum of emitted particles tends (i.e., is more likely) to carry away angular momentum and electric charge. Because of the characteristics of the spectrum, a Kerr-Newman black hole will quickly (in comparison with time scales of interest) become a Schwarzschild black hole. Therefore, for simplicity, our analysis will be focused on a Schwarzschild black hole.
@@
As per equation (4), the Hawking temperature of a Schwarzschild black hole is inversely proportional to its mass, which means that the typical energy of created particles increases as a Schwarzschild black hole evaporates. In particular, this relation between particle energy spectrum and black hole mass ensures that created particles will have a small influence on the spacetime metric for most of the evaporation process, so that neglecting additional backreaction effects will be a good approximation when $r_{s}\gg\ell_{p}$ (i.e., far from the Planck scale). In this sense, details of backreaction effects will pose little influence over the event horizon and the exterior region. Consequently, the spacetime geometry can be described by a sequence of quasi-static processes in which the black hole radius decreases slowly, with the process of energy loss being approximated by Stefan’s law. This argument leads one to the conclusion that a black hole should completely evaporate, i.e., radiate away its mass, in a finite amount of time (as “measured” by Cauchy hypersurfaces). Evidently, at some point in time $r_{s}\sim\ell_{p}$, hence, complete evaporation is only a valid prediction if one considers that no deviations from SG occur at that scale. For the moment, let us consider that to be the case and analyze the consequences of a complete evaporation process.
The conformal diagram of a Schwarzschild black hole that resulted from a spherically symmetric collapse and evaporated completely is illustrated in fig. 12. Notably, the spacetime fails to be globally hyperbolic.
Figure 12: Conformal diagram of a Schwarzschild black hole that formed from a spherically symmetric distribution of energy and evaporated completely.
@@
This can be verified by the fact that some null curves do not intersect $\Sigma_{2}$. More precisely, although after complete evaporation events at $r=0$ are again causally connected to $\mathscr{I}^{+}$, there are null curves that never intersect $\Sigma_{2}$ because they reach the singularity [^9] Notice that the event of complete evaporation corresponds to a singularity that is causally connected to $\mathscr{I}^{+}$. Thus, null curves emanating from it do not intersect any section such as $\Sigma_{0}$ or $\Sigma_{1}$.. Thus, physical conditions at $\Sigma_{2}$ will not suffice to determine those at $\Sigma_{0}$ or $\Sigma_{1}$. Fundamentally, information loss can be attributed to the fact that some null curves reach the singularity. Indeed, information loss is also evident from the fact that the field state at a spacelike section such as $\Sigma_{2}$ (“after” complete evaporation) is not sufficient to determine the complete field state at either $\Sigma_{0}$ or $\Sigma_{1}$. In particular, considering the same color scheme of fig. 11, at $\Sigma_{0}$ one has an initial pure Hadamard state. At $\Sigma_{1}$, the total field state is still pure, but the field state outside the black hole is a mixed one due to entanglement with the black hole interior. Consequently, after complete evaporation, the field state at any spacelike section such as $\Sigma_{2}$ will be a mixed one. In essence, there is still entanglement with black hole interior, even though it “no longer” exists. It is in this sense that black holes will result in a dynamical evolution from a pure state to a mixed one (i.e., a non-unitary dynamical evolution).
@@
The generality of this analysis follows from the arguments presented in Secs. II and III. Namely, suppose one starts with an isolated energy distribution described by a pure Hadamard state. If this distribution is massive enough, gravitational effects will not be sustained, and a black hole will form. Following the emission of gravitational waves, the black hole will eventually be described to a good order of approximation by a Kerr-Newman configuration. The Hawking effect together with the qualitative description of backreaction means that the Kerr-Newman black hole will (quickly) result in a Schwarzschild black hole. One is then led to the same line of reasoning presented above, where information about the field state after complete evaporation does not suffice to determine the initial state.
@@
Now, the expectation of information conservation is only justified in a globally hyperbolic spacetime, where information is carried from one Cauchy surface to another. In fact, QFT predicts that information on a Cauchy surface will not in general be completely deposited in a future region that fails to be a Cauchy surface. This means that even in a globally hyperbolic spacetime one can still find examples of “information loss”. For example, in Minkowski spacetime one can consider the field evolution from a Cauchy surface, $\Sigma_{0}$, to an asymptotically null hyperboloid, $\Sigma_{1}$ (see fig. 13) Wald1994 .
@@
Because there are null curves that do not intersect $\Sigma_{1}$, a pure state at $\Sigma_{0}$ would necessarily evolve to a mixed one at $\Sigma_{1}$. As such, dynamical evolution from $\Sigma_{0}$ to $\Sigma_{1}$ would not be unitary as it does not constitute a closed system (i.e., an isolated system). Evidently, this is not at odds with the postulates of QM, which requires unitary evolutions only for closed systems.
@@ IV.2 Misconceptions about alternatives to information loss
It has been extensively advocated in the literature that such pure-to-mixed-state evolution implies a violation of physical principles and conjectures regarding black holes (e.g., that black holes are ordinary thermodynamic systems Almheiri2021 ), and therefore, black hole evaporation must be a unitary process. The incorrect line of reasoning that leads to these interpretations views information loss as paradoxical, and refers to the state of affairs surrounding the product of black hole evaporation as the black hole information paradox.
Notably, if one proposes a different description of black hole evaporation that leads to information conservation, it is imperative to preserve semiclassical physics far from the Planck scale. More precisely, a proposal that seeks to change semiclassical predictions at arbitrary regimes is inherently contradictory. Examples of such contradictory proposals are those that support the view that black holes never form (e.g., Mathur2005 ) and those arguing that Hawking radiation should strongly deviate from thermality (e.g., Almheiri2021 ). Notice that formation of black holes and the derivation of the predominant thermal character of the Hawking radiation are robust results that do not rely on specific regimes. For instance, consider the curvature scales at the event horizon of a Schwarzschild black hole, which are proportional to $r_{s}^{-2}$. From eq. 2, in order to avoid the formation of any Schwarzschild black hole, one would have to find deviations from SG predictions in arbitrary curvature scales. The exact same argument applies to deviations from thermality by Hawking radiation. That is, a strong deviation from thermality at arbitrary regimes could only be accounted for by strong deviations from the semiclassical picture far from the Planck scale. However, QFT and GR are well-tested at low curvature regimes (e.g., Solar system-like curvature settings), and, because no deviations are found, relying on arguments that would require such descriptions to fail at arbitrary regimes is contradictory. On the other hand, nothing can be said about proposals that change the product of the evaporation process only at the Planck scale, neither in favor nor against them.
@@ V Conclusions
In this paper, we reviewed the classical and semiclassical description of black holes and presented the conjectures and hypotheses that lead to the conclusion of information loss. We show that the physical non-unitary dynamical evolution that follows from black hole complete evaporation can be interpreted as a consequence of the singularity “opening” the system. Consequently, the conclusion of information loss in SG is not paradoxical. Additionally, we discussed the contradictory nature of proposals that suggest deviations from SG at arbitrary scales.
While acknowledging that a theory of quantum gravity is required to make precise predictions about the outcome of black hole evaporation, particularly regarding backreaction effects and Planck scale physics, we emphasize that the hypotheses, conjectures, and approximations of SG are well justified. Nonetheless, it remains a compelling open issue how black hole evaporation will be addressed within an adequate theory of quantum gravity, and what novel physics it could provide.
@@ Acknowledgments
The author thanks Esmerindo de Sousa Bernardes for his support during the early stages of this work, Daniel A. T. Vanzella for extensive support and for comments on a draft, and George E. A. Matsas for many conversations on black hole physics and support during the later stages of this work. The author also thanks two anonymous referees for detailed and helpful comments. This study was financed by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001.

### v4

@@ IV Time evolution of semiclassical black holes
In this section, we show how the semiclassical properties of black holes discussed so far imply that information is lost. We will show why this conclusion is in agreement with semiclassical arguments, and does not earn the adjective “paradoxical”. Additionally, we will present misconceptions about proposals that seek to change the conclusion regarding information loss.
@@ IV.1 Information loss
Although calculating backreaction effects is not a trivial task for four-dimensional spacetimes, one expects that the main consequence of the Hawking radiation is to reduce the black hole parameters at a given time, i.e., its mass, angular momentum, and electric charge. This results in an effective evaporation process. This qualitative description of a possible backreaction effect is corroborated by physical arguments. More precisely, the energy reduction is merely a consequence of the fact that the positive energy flux to $\mathscr{I}^{+}$ implies that there must exist a negative energy flux going into the black hole, while the reduction of the other parameters is because the spectrum of emitted particles tends (i.e., is more likely) to carry away angular momentum and electric charge. Because of the spectrum characteristics, a Kerr-Newman black hole will quickly (in comparison with time scales of interest) become a Schwarzschild black hole. Therefore, for simplicity, our analysis will be focused on a Schwarzschild black hole.
@@
As per eqs. 4 and 5, the Hawking temperature of a Schwarzschild black hole is inversely proportional to its mass, which means that the typical energy of created particles increases as a Schwarzschild black hole evaporates. In particular, this relation between particle energy spectrum and black hole mass ensures that created particles will have a small influence on the spacetime metric for most of the evaporation process, so that neglecting additional backreaction effects will be a good approximation when $r_{s}\gg\ell_{p}$ (i.e., far from the Planck scale). In this sense, details of backreaction effects will have little influence over the event horizon and the exterior region. Consequently, the spacetime geometry can be described by a sequence of quasi-static processes in which the black hole radius decreases slowly, with the process of energy loss being approximated by Stefan’s law. This argument leads one to the conclusion that a black hole should completely evaporate, i.e., radiate away its mass, in a finite amount of time (as “measured” by Cauchy hypersurfaces). Evidently, at some point in time $r_{s}\sim\ell_{p}$, hence, complete evaporation is only a valid prediction if one considers that no deviations from SG occur at that scale. For the moment, let us consider that to be the case and analyze the consequences of a complete evaporation process.
@@
The conformal diagram of a Schwarzschild black hole that results from a spherically symmetric collapse and completely evaporates is illustrated in fig. 12. Notably, the spacetime fails to be globally hyperbolic. This can be verified by the fact that some null curves do not intersect $\Sigma_{2}$. More precisely, although after complete evaporation events at $r=0$ are again causally connected to $\mathscr{I}^{+}$, there are null curves that never intersect $\Sigma_{2}$ because they reach the singularity [^11] Note that the event of complete evaporation corresponds to a singularity that is causally connected to $\mathscr{I}^{+}$. Thus, null curves emanating from it do not intersect any section such as $\Sigma_{0}$ or $\Sigma_{1}$.. Hence, physical conditions at $\Sigma_{2}$ will not suffice to determine those at $\Sigma_{0}$ or $\Sigma_{1}$. Fundamentally, information loss can be attributed to the fact that some null curves reach the singularity. Indeed, information loss is also evident from the fact that the field state at a spacelike section such as $\Sigma_{2}$ (“after” complete evaporation) is not sufficient to determine the complete field state at either $\Sigma_{0}$ or $\Sigma_{1}$. In particular, considering the same color scheme of fig. 11, at $\Sigma_{0}$ one has an initial pure Hadamard state. At $\Sigma_{1}$, entanglement due to the Hadamard state means that the field state outside the black hole will be a mixed one. Consequently, after complete evaporation, the field state at any spacelike section such as $\Sigma_{2}$ will also be a mixed one.
Figure 12: Conformal diagram of a Schwarzschild black hole that forms from a spherically symmetric energy distribution and evaporates completely.
@@
The generality of this analysis follows from the arguments presented in secs. II and III. Namely, suppose one starts with an isolated energy distribution described by a pure Hadamard state. If this distribution is sufficiently massive, gravitational effects will not be sustained, and a black hole will form. Following the emission of gravitational waves, the black hole exterior will eventually be well approximated by a Kerr-Newman metric. The Hawking effect, together with the qualitative description of backreaction, indicates that the Kerr-Newman black hole will (quickly) result in a Schwarzschild black hole. One is then led to the same line of reasoning presented above, where information about the field state after complete evaporation does not suffice to determine the initial state.
@@
Now, the expectation of information conservation is only justified in a globally hyperbolic spacetime, where information is carried from one Cauchy surface to another. In fact, QFT predicts that information on a Cauchy surface will not, in general, be completely deposited in a future region that fails to be a Cauchy surface. This means that even in a globally hyperbolic spacetime, one can still find examples of “information loss” (in the sense that a pure state evolves to a mixed one). For instance, in Minkowski spacetime, one can consider the field evolution from a Cauchy surface, $\Sigma_{0}$, to an asymptotically null hyperboloid, $\Sigma_{1}$ (see fig. 13) Wald1994 . Because there are null curves that do not intersect $\Sigma_{1}$, a pure state at $\Sigma_{0}$ would necessarily evolve to a mixed one at $\Sigma_{1}$. As such, dynamical evolution from $\Sigma_{0}$ to $\Sigma_{1}$ would not be unitary as it does not constitute a closed system (i.e., an isolated system). Evidently, this is not at odds with the postulates of QM, which require unitary evolutions only for closed systems.
@@ IV.2 Misconceptions about alternatives to information loss
It has been extensively argued in the literature that a pure-to-mixed-state evolution violates physical principles and contradicts conjectures about black holes (e.g., that black holes are ordinary thermodynamic systems Almheiri2021 ), and therefore black hole evaporation must be a unitary process. The incorrect line of reasoning that leads to these interpretations views information loss as paradoxical, and refers to the state of affairs surrounding the product of black hole evaporation as the black hole information paradox.
Notably, if one proposes a different description of black hole evaporation that leads to information conservation, it is imperative to preserve semiclassical physics far from the Planck scale. More precisely, a proposal that seeks to change semiclassical predictions at arbitrary regimes is inherently contradictory. Examples of such contradictory proposals are those that support the view that black holes never form (e.g., Mathur2005 ) and those arguing that Hawking radiation should strongly deviate from thermality (e.g., Almheiri2021 ). Note that formation of black holes and the derivation of the predominant thermal character of the Hawking radiation are robust results that do not rely on specific regimes. For instance, consider the curvature scales at the event horizon of a Schwarzschild black hole, which are proportional to $r_{s}^{-2}$. From eq. 2, to avoid the formation of any Schwarzschild black hole, one would have to find deviations from SG predictions in arbitrary curvature scales. The exact same argument applies to deviations from thermality by Hawking radiation. That is, a strong deviation from thermality at arbitrary regimes could only be accounted for by strong deviations from the semiclassical picture far from the Planck scale. However, QFT and GR are well-tested at low curvature regimes (e.g., solar system-like curvature settings), and, because no deviations are found, relying on arguments that would require such descriptions to fail at arbitrary regimes is contradictory. On the other hand, nothing can be said about proposals that change the product of the evaporation process only at the Planck scale, neither in favor nor against them.
@@ V Conclusions
In this paper, we reviewed the classical and semiclassical descriptions of black holes and presented the conjectures and hypotheses that lead to the conclusion of information loss. We show that the physical non-unitary dynamical evolution that follows from black hole complete evaporation can be interpreted as a consequence of the singularity “opening” the system. Consequently, the conclusion of information loss in SG is not paradoxical. Additionally, we discussed the contradictory nature of proposals that suggest deviations from SG at arbitrary scales.
While acknowledging that a theory of quantum gravity is required to make precise predictions about black hole evaporation, particularly regarding backreaction effects and Planck-scale physics, we emphasize that the hypotheses, conjectures, and approximations of SG are well justified. Nonetheless, it remains a compelling open issue how black hole evaporation will be addressed within an adequate theory of quantum gravity, and what novel physics it could provide.
@@ Acknowledgments
The author thanks Esmerindo de Sousa Bernardes for his support during the early stages of this work, Daniel A. T. Vanzella for extensive support and comments on a draft, and George E. A. Matsas for many conversations on black hole physics and support during the later stages of this work. The author also thanks two anonymous referees for their detailed and helpful comments. This study was financed by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001, and by the Sao Paulo Research Foundation (FAPESP), Brazil. Process Number 2024/19057-7.